can somw one help??? thank you
90. CD rollover
Ronnie invested p dollars in a 2 year CD with the annual rate of return of r. After the CD rolled over three times, its value was p (1+r) ^2) ^3. Which law of exponents can be used to simplify the expression? Simplify it.

Answer by stanbon(75887)

p (1+r) ^2) ^3. Which law of exponents can be used to simplify the expression? Simplify it.
Comment: There is an "open parenthesis" missing.
I'll assume the expression is
(p(1+r)^2)^3
=p^3 (1+r)^6
Because a^m^n=a^(mn)
Cheers,
Stan H.